OpenStudy (anonymous):

A rectangular yard is surronded by a fence that is 4 feet longer than it is wide. if the perimeter of the fence is 424 feet what are the dimensions of the yard it encloses?

OpenStudy (anonymous):

P=2L+2W for rectangles. The fence is 4 feet longer than it is wide. This translates into L=W+4. Substitute this L value along with your perimeter value and solve. P=2L+2W 424=2(W+4)+2W 424=2W+8+2W 424=4W+8 416=4W ***104=W*** We know that L=W+4, so let's plug in our new W value. L=W+4 L=104+4 ***L=108*** The yard is 108 ft by 104 ft.

OpenStudy (anonymous):

thanks!!! can you help with few more?